[figure: 9.7 cm]

ANOMALOUS BEHAVIOR OF MAGNETOSTRICTION AT METAMAGNETIC PHASE TRANSITIONS IN MODEL ANTIFERROMAGNET DyAlO_{3}

I.B. Krynetskii^{a}, V.M. Matveev^{b}, V.V. Matveev^{b} aMoscow State University, 119899, Moscow, Russia

bState Research Institute for Physical Problems, 103460, Zelenograd, Moscow, Russia

Unusual static and dynamic behavior of low-temperature (1.5-4.2 K) magnetostriction of Ising Dy^{3+} ions in dipolar antiferromagnet DyAlO_{3} at metamagnetic phase transitions induced by external magnetic fields along a- and b-axes (up to 5 T) was observed. The magnetostriction curves was found to be strongly dependent on the rate of rise and direction of change of the field showing odd-shaped hysteresis and dynamic enhancement of magnetostriction jumps. A ten-fold increase in the rate of field rise resulted in approximatively four-fold increase in magnetostriction jumps. It is interesting that the magnetostriction jumps had a maximum at T > T_{N} when magnetic sublattices were not saturated fully and decreased practically to zero with further decreasing temperature. A theoretical model for explanation of such behavior taking into account nonuniform magnetic states of the sample due to the staggered field and ”magnetocaloric focussing” of magnetostriction anomalies was proposed. The work was supported by RFBR (grant 97-02-17643, 98-02-16469).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : I.Krynetskii

Address for correspondence : Department of Physics Moscow State University 119899, Moscow, Russia

Email address : krynets@plms.phys.msu.su